how to would i solve
2m+n=13
m-2n=14

Answer by mananth(16949)

2m+n=13 .............1
m-2n=14 .............2

multiply (1)by 2
Multiply (2) by 1
4m+2n=26
m-2n=14
Add the two equations
5m=40
/5
m=8
plug value of m in (1)
2m+n=13
16+n=13
n=13-16
n=-3